OpenStudy (anonymous):

There are 20 students in a club, 12 boys and 8 girls. If 5 members of the cub are chosen at random, what is the probability that in the group chosen there are exactly 2 boys?

OpenStudy (anonymous):

Given: 20 students in total 12 boys 8 girls • If there is no any restrictions, there are \(\sf  _{20}C_{5}\) ways that they can choose the members. • 5 members are needed and there should be EXACTLY two boys in the club. This means we need three girls and 2 boys to complete the club. So \(\sf   _{8}C_{3} \times  _{12}C_{2}\) will be the number of ways that exactly two boys will be included in the team. Now, probability will be \(\sf \Large \frac{\sf   _{8}C_{3} \times  _{12}C_{2}}{  _{20}C_{5}}\)

OpenStudy (anonymous):

btw, welcome to openstudy @cjsyki

OpenStudy (anonymous):

thanks!